# AdaDec: Uncertainty-Guided Adaptive Decoding for LLM-based Code Generation

### v1

@@ Abstract
Code generation with large language models (LLMs) is highly sensitive to token selection during decoding, particularly at uncertain decision points that influence program logic. While standard strategies like greedy and beam search treat all tokens uniformly, they overlook code-specific uncertainty patterns, leading to suboptimal performance. This paper presents an empirical study revealing that many generation errors stem from ranking mistakes at high-uncertainty steps, where the correct token is present but not top-ranked.
Motivated by these findings, we propose AdaDec, an uncertainty-guided adaptive decoding framework that integrates a token-level pause-then-rerank mechanism driven by token uncertainty (Shannon entropy). AdaDec learns model-specific uncertainty thresholds and applies a lookahead-based reranking strategy when uncertainty is high. Experiments on HumanEval and MBPP benchmarks show that AdaDec improves Pass1 accuracy by up to 15.5% over greedy decoding, outperforms or matches beam search, and reduces computational cost and latency through efficient, selective pausing. Our results highlight the promise of uncertainty-aware adaptive decoding for improving the reliability and efficiency of LLM-based code generation.

### v2

# Towards Better Code Generation: Adaptive Decoding with Uncertainty Guidance

@@ Abstract
Code generation using large language models (LLMs) is highly sensitive to the choice of tokens during decoding, especially at points of uncertainty that critically affect the generated program’s logic. Conventional decoding methods such as greedy search and beam search apply uniform treatment to all tokens, neglecting the unique uncertainty characteristics inherent in code generation, which can result in suboptimal outputs. In this work, we conduct an empirical analysis demonstrating that a significant portion of generation errors arises from incorrect token ranking at high-uncertainty steps—where the ground truth token exists in the candidate set but fails to be ranked first.
Inspired by this insight, we introduce AdaDec, an adaptive decoding framework guided by token-level uncertainty quantified via Shannon entropy. AdaDec dynamically learns uncertainty thresholds tailored to each model and employs a pause-then-rerank mechanism with lookahead when the uncertainty surpasses these thresholds. Evaluation on the HumanEval and MBPP benchmarks reveals that AdaDec achieves up to a 15.5% improvement in Pass@1 accuracy compared to greedy decoding, matches or outperforms traditional beam search, and reduces both computational overhead and latency through targeted, selective pausing. Our findings suggest that uncertainty-aware adaptive decoding holds considerable potential for enhancing both the reliability and efficiency of code generation with LLMs.